What are your preferred time(s) for the Thursday discussion section?

10:30 - 11:20 11:30 - 12:20 12:30 - 1:20

What is your preferred date(s) for exam 2? What day(s) do you NOT want to have exam 2? On a scale of 1-10, is the class too fast/too slow?

too slow = 1, too fast = 10

How many hours per week do you spend on this course (include class attendance)? How do you find my use of the overhead projector/computer? Other comments/suggestions: